OpenStudy (anonymous):

Find all real solutions of the equation by completing the square. 4x^2 + 5x − 4 = 0

OpenStudy (blurbendy):

x = 1/8 (-5 - Sqrt[89]) x = 1/8(Sqrt[89] - 5)

OpenStudy (anonymous):

Thats actually the answer i got but im wondering why it says it is wrong on my online homework...? Thanks thou !!

OpenStudy (anonymous):

In your opinion do you think putting the answer as a decimal is wrong?

OpenStudy (anonymous):

because when i put the answer the way you wrote it , it said it was correct

OpenStudy (blurbendy):

you put it in as a decimal? It depends on what the question wanted. If it wanted the answer in the exact form then you would enter it like how I did. If it asked for a numerical approximation, or a decimal, then you would use a decimal. If it says neither, then just enter it in the exact form.

OpenStudy (anonymous):

Alright thank you :) yeah it wasn't specified but that's what i don't like about this online stuff ... sometimes it takes the decimal and sometimes doesn't thanks a lot though :)

OpenStudy (blurbendy):

no problem! if you have any other questions just ask